Point charges of 21.0 µC and 40.0 µC are placed 0.500 m apart. (a) At what point (in m) along the line connecting them is the electric field zero? m (from the smaller charge) (b) What is the magnitude (in N/C) and direction of the electric field halfway between them? magnitude N/C direction
Answer #1

here,

a) let the distance is x from 21 uC

k * q1/x^2 = k * q2/(d - x)^2

21/x^2 = 40/(0.50 - x)^2

solving for x

x = 0.21 m

the distance at which electric field is zero is 0.21 m

b) at the midpoint

d = x/2 = 0.50/2 = 0.25 m

Electric field = 9 *10^9 * 10^-6 * (40 - 21)/0.25^2

Electric field = 2.74 *10^6 N/C

the electric field is 2.74 *106 N/C

the direction of field is towards smaller charge